# Quantum metrology of low frequency electromagnetic modes with frequency upconverters

### v2

@@ VI RQU Upconversion Demonstration
The high extinction ratio proves the viability of phase-sensitive upconversion, although it does not constitute a full backaction-evading measurement, which will require a high-Q resonant circuit on the input of the upconverter. For this signal frequency and this upconverter, the spurious backaction terms $n_{bad}$ would have limited the degree of backaction evasion to less than 10dB, in any case. However, reducing the microwave loss rate $\kappa$ by increasing the microwave quality factor can reduce the spurious backaction, in principle arbitrarily.
@@ VII Conclusions
The RQU a tool for quantum metrology of low-frequency electromagnetic modes. We have shown that this interaction allows the RQU to act as a quantum-limited op-amp with tunable noise impedance. Using the RQU as a replacement for dc SQUIDs enables amplification at the SQL and in situ tunable noise impedance. Furthermore, the RQU can achieve an uncoupled “energy sensitivity” $(1/2)L_{\rm in}S_{\rm I}$ significantly below $\hbar$ when coupled to an untuned input inductor. Such performance is not possible with a dc SQUID in which $L_{\rm in}$ is linked to $S_{\rm I}$ Voss (1981). These functions are useful in a variety of magnetometry applications.

### v3

@@ VI RQU Upconversion Demonstration
The high extinction ratio proves the viability of phase-sensitive upconversion, although it does not constitute a full backaction-evading measurement, which will require a high-Q resonant circuit on the input of the upconverter. For this signal frequency and this upconverter, the spurious backaction terms $n_{\rm{bad}}$ would have limited the degree of backaction evasion to less than 10dB, in any case. However, reducing the microwave loss rate $\kappa$ by increasing the microwave quality factor can reduce the spurious backaction, in principle arbitrarily.
@@ VII Conclusions
We have demonstrated that the RQU is a valuable tool for quantum metrology of low-frequency electromagnetic modes. We have shown that this interaction allows the RQU to act as a quantum-limited op-amp with tunable noise impedance. Using the RQU as a replacement for dc SQUIDs enables amplification at the SQL and in situ tunable noise impedance. Furthermore, the RQU can achieve an uncoupled “energy sensitivity” $(1/2)L_{\rm in}S_{\rm I}$ significantly below $\hbar$ when coupled to an untuned input inductor. Such performance is not possible with a dc SQUID in which $L_{\rm in}$ is linked to $S_{\rm I}$ Voss, (1981). Since it is capable of operating at the SQL, with tunable noise impedance, and achieving $\hbar$ uncoupled energy sensitivity the RQU exceeds the capability of the current state-of-the-art, dc SQUIDs. The capabilities provided by the RQU can be useful in a variety of magnetometry applications.